# Zeroth-order log-concave sampling

### v1

@@ 1.2 Technical overview
1.3 Related work
We survey the literature on applications of zeroth-order log-concave sampling and log-concave sampling in other settings (e.g., given access to a different set of oracles).
More on uniform sampling.
[…]
As mentioned earlier, Lovász and Simonovits [LS93] used the $\mathsf{Ball\ walk}$ to develop an indirect approach with improved complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$. It was achieved by refining their earlier conductance framework and localization machinery; they proved that any log-concave distribution $\pi$ supported on a convex body of diameter $D$ also satisfies $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$.
Recently, $\mathsf{Coordinate\ Hit\text{-}and\text{-}Run}$ ($\mathsf{CHR}$), a special variant of $\mathsf{Hit\text{-}and\text{-}Run}$, has been shown to mix in polynomial time with respect to the dimension and ($\ell_{2}$ or $\ell_{\infty}$) diameter [NS22, LV23, FV23], and with a logarithmic dependence on the initial warmness [NRS25]. This thus results in $\mathsf{CHR}:\mathcal{R}_{2}\to\mathcal{R}_{2}$.
KLS conjecture.
The Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] posits that the Cheeger constant of any log-concave probability measure is bounded below by $\Omega(\lVert\operatorname{cov}\pi\rVert^{-1/2})$. Starting from $C_{\mathsf{Ch}}^{2}\gtrsim\operatorname{tr}^{-1}(\operatorname{cov})$, subsequent work [Eld13, LV24, Che21, KL22, Kla23]—notably through Eldan’s stochastic localization—has improved it to $C_{\mathsf{Ch}}^{2}\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [Kla23]. However, the improved Cheeger constant in this setting does not translate to a better Cheeger constant with respect to the cross-cut ratio.
Rounding.
A notable application is the rounding of convex bodies, or more generally, log-concave functions. It aims to make the geometry of the input object more regular—typically by applying an affine transformation that brings it into a nearly-isotropic position. This step is often a necessary preprocessing phase for achieving a fully polynomial-time algorithm, since worst-case shapes (e.g., highly skewed bodies) can cause sampling algorithms to suffer from poor mixing.
Earlier approaches to rounding relied on deterministic methods based on the ellipsoid algorithm [DFK91, LS90], which guarantees $B_{1}(0)\subset\mathcal{K}\subset B_{d^{3/2}}(0)$ after a suitable affine transformation. To improve the sandwiching ratio, randomized rounding methods were pioneered in [LS93] and further developed in [KLS97, LV06c, JLLV21, JLLV24]. These methods use sampling to estimate the covariance matrix and then apply affine transformations to round the body. The best known complexity for isotropic rounding of a convex body is $\widetilde{O}(d^{3.5}\operatorname{polylog}R)$ [JLLV24], with the analysis streamlined in [KZ24]. These rounding techniques have also been extended to log-concave distributions [LV06a, KV25b], while preserving the query complexity.
Volume computation and integration.
This framework initiated a long line of research focused on improving the mixing time, oracle complexity, and practical implementability of such algorithms [Lov90, AK91, LS90, LS93, KLS95, KLS97, LV06b, CV15, CV18, JLLV21]. The current fastest algorithm for isotropic convex bodies is due to [CV18] with complexity $\widetilde{O}(d^{3}/\varepsilon^{2})$. A natural generalization of volume computation is the integration of a log-concave function. This general problem was studied in [LV06a, KV25b], with the same complexity $\widetilde{O}(d^{3}/\varepsilon^{2})$. See [KV25b, §5.2] for a streamlined analysis.
Sampling with additional assumptions.
For constrained sampling problems—such as the convex-body sampling—there exist approaches beyond general-purpose MCMC algorithms like the $\mathsf{Ball\ walk}$, $\mathsf{Hit\text{-}and\text{-}Run}$, and $\mathsf{PS}_{\textup{unif}}$. These alternatives often leverage additional geometric structure of the constraint set $\mathcal{K}$, assume additional regularities of target distributions (e.g., smoothness), or assume access to stronger oracles, such as projection or first-order oracles.
(1) Sampling with local geometry. This approach assumes the existence of a tractable self-concordant barrier $\phi$ for the constraint $\mathcal{K}$. At each step, a sampler draws from an anisotropic Gaussian distribution whose covariance is pertinent to the local geometry encoded by $\nabla^{2}\phi$. This leads to an affine-invariant sampler that naturally adapts to the shape of $\mathcal{K}$, without requiring a separate rounding procedure. The most notable example is the $\mathsf{Dikin\ walk}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ [KN12], which, when used with a log-barrier for a polytope with $m$ linear constraints, achieves a mixing time of $O(md\log\nicefrac{{M_{2}}}{{\varepsilon}})$, and its analysis was streamlined in [SV16].
This barrier-based sampler has been generalized and improved through more efficient barrier functions [Nar16, CDWY18, GN23]. A unifying framework incorporating these developments—as well as procedures for warm-start generation—was presented in [KV24]. Moreover, complexity analysis beyond worst case was explored in [JC24], and alternative non-Euclidean schemes using log-Laplace transformations was proposed in [GLL+23].
(2) Sampling with first-order oracle. When a first-order oracle is available, faster non-Euclidean samplers become feasible. These include Riemannian Hamiltonian Monte Carlo  [GC11, LV18, GKV24], which uses Hamiltonian dynamics on the Hessian manifold with metric $\nabla^{2}\phi$, and Riemannian Langevin Algorithm [GC11, LE23], which extends Langevin dynamics to a Riemannian manifold. Other approaches include the Mirror Langevin Algorithm [ZPFP20, Jia21, AC21, LTVW22, SWW24], which use mirror maps to simulate Langevin-type updates in dual space, and the Preconditioned Langevin Algorithm [SWW25].
(3) Sampling with projection oracle. In this setting, access to a projection oracle—computing the projection of a point onto $\mathcal{K}$—enables other sampling methods: unconstrained steps (e.g., Euler–Maruyama discretization of the Langevin diffusion) are followed by projections back into the feasible set [BEL18, Leh23]. Alternatively, soft-constraint approaches have also been proposed, in which the target distribution is modified with penalty terms that discourage constraint violations while remaining close to the original distribution. Such penalized schemes have been studied in [BDMP17, GHZ24, Yu24].
(4) Zeroth-order sampling for well-conditioned distributions. Those three general-purpose samplers—the $\mathsf{Ball\ walk}$, $\mathsf{Hit\text{-}and\text{-}Run}$, and $\mathsf{Proximal\ sampler}$—have also been studied by [ALPW24], [ALZ24], and [LST21, CCSW22] in the unconstrained well-conditioned setting with access to an evaluation oracle. These samplers achieve the mixing times of $\widetilde{O}(\kappa d\log\nicefrac{{1}}{{\varepsilon}})$ under various metrics (e.g., $\mathcal{W}_{2}$, $\mathsf{KL}$, $\chi^{2}$).
Real-world applications and impacts.
Sampling research is not only of theoretical interest but also has far-reaching implications for real-world applications. A prominent application arises in systems biology, where uniform sampling from a polytope can be used to analyze the efficiency of human metabolic networks [T+13, HCT+17]. Advances in sampling techniques have enabled more scalable biological inferences. For instance, the development of non-Euclidean samplers has led to improved mixing times and computational efficiency [YK17, CF21, CK21, KLSV22, KLSV23, SC24]. Moreover, practical implementations of volume computation for convex bodies have been developed in [CV16], bridging the gap between theoretical developments and real-world usability. Collectively, these advancements demonstrate how sampling research continues to yield practical tools with deep and lasting impact across science and engineering disciplines.
@@ 1.4 Preliminaries
We call a function $f:\mathbb{R}^{d}\to[0,\infty)$ log-concave if $-\log f$ is convex in $\mathbb{R}^{d}$, and a probability measure $\pi$ (or distribution) log-concave if it has a log-concave density function with respect to the Lebesgue measure. We abuse notation by using the same symbol for a distribution and density. We assume that any log-concave distributions considered in this work are non-degenerate; otherwise, we could simply work on an affine subspace on which degenerate distributions are supported. For $t\geq 0$, $\pi$ is called $t$-strongly log-concave if $-\log\pi$ is $t$-strongly convex (i.e., $-\log\pi-\frac{t}{2}\,|\cdot|^{2}$ is convex). Clearly, log-concavity is preserved under multiplication, and a classical result by Prékopa and Leindler ensures that the convolution also preserves log-concavity. A distribution is called isotropic if its barycenter is at the origin and has the identity covariance matrix. Note that as the log-concave distributions decay exponentially fast at infinity, they have finite moments of all orders.
@@
For $\lambda=q/(2q-1)$, we have $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\frac{2q}{2q-1}\,\mathcal{R}_{2q-1}(\mu\mathbin{\|}\nu)+\mathcal{R}_{2q}(\nu\mathbin{\|}\pi)$. We refer readers to [vEH14, Mir17] for more properties of the Rényi divergence.
2 Uniform sampling from convex bodies
We begin by describing our approach for the general proximal sampler ($\mathsf{PS}$) and then focus specifically on $\mathsf{PS}_{\textup{unif}}$ in §2.2. $\mathsf{PS}$ with step size $h$ alternates between two steps: starting with $k=0$ and $x_{0}\sim\pi_{0}^{X}$, (i) sample $y_{k+1}\sim\pi^{Y|X=x_{k}}$ and (ii) sample $x_{k+1}\sim\pi^{X|Y=y_{k+1}}$. Precisely, let $P_{h}$ and $Q_{h}$ be the transition kernel of the forward and backward step respectively, defined as
@@ 2 Uniform sampling from convex bodies
Assume that there exists a suitable proposal distribution $\mu_{y}$ such that $(\mu_{y}\,|\,\text{accepted})=Q_{h}(y,\cdot)$. In case of uniform sampling, we use $\mu_{y}=\mathcal{N}(y,hI_{d})$ and accept the proposal if it lies inside of $\mathcal{K}$.
Given this setup, we implement the backward step using rejection sampling with proposal $\mu_{y_{k+1}}$. If this rejection loop exceeds $\tau$ trials, we discard $y_{k+1}$ and return to the forward step using the same $X=x_{k}$. We group each forward–backward pair (with rejection sampling up to $\tau$ retries) into a stage. For instance, if all $\tau$ retries fail in stage $1$, then we begin stage $2$ by drawing a new $\bar{y}_{k+1}\sim\pi^{Y|X=x_{k}}$ and sampling from $x_{k+1}\sim\pi^{X|Y=\bar{y}_{k+1}}$. Stages are repeated until one iteration completes successfully.
We now analyze the query complexity of $\mathsf{PS}$ for obtaining a sample whose law is $\varepsilon$-close to the target $\pi^{X}$ in a chosen metric (Theorem 1.1). To this end, we decouple the analysis into two components: (1) a mixing analysis and (2) a query complexity of the backward step. The first part determines how many iterations $N$ are required for the sample $X_{N}$ to be $\varepsilon$-close to $\pi^{X}$. The second one determines how many queries $\mathsf{PS}$ uses in each iteration. Therefore, the total query complexity of the algorithm is the product of these two components.

### v2

@@ 1.1.2 Improved warm-start generation via annealing
1.1.3 Logconcave sampling under an evaluation oracle
Our uniform-sampling results extend naturally to general logconcave distributions $\pi\propto e^{-V}$, given access to an evaluation oracle for a convex function $V:\mathbb{R}^{d}\to\mathbb{R}\cup\{\infty\}$. In this setting, the regularity condition $B_{1}\subset\mathcal{K}$ translates to a condition on a level set of $V$, namely, that the ground set $\mathsf{L}_{\pi,g}:=\{x\in\mathbb{R}^{d}:V(x)-\min V\leq 10d\}$ contains a unit ball. This generalization is justified by the fact that $\mathsf{L}_{\pi,g}$ takes up a constant fraction of the $\pi$-measure (see [KV25b, Lemma 2.6]). We emphasize that any logconcave distribution satisfies this condition after rescaling.
The logconcave sampling problem has been studied as an algorithmic problem since Applegate and Kannan [AK91], who used the $\mathsf{Grid}\text{ }\mathsf{walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$, establishing polynomial complexity in the dimension under the assumption of Lipschitzness of the function over the support. Later, Lovász and Vempala [LV07] generalized classical algorithms such as the $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$ to this setting by incorporating a Metropolis filter. They showed that after a suitable pre-processing of $\pi$, $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{\infty}\to\mathsf{TV}$ with query complexity $\widetilde{O}(\varepsilon^{-4}d^{2}R^{2}M_{\infty}^{4}\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$. Furthermore, [LV06a] showed that the complexity of $\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ can be improved to $\widetilde{O}(d^{2}R^{2}\operatorname{polylog}\nicefrac{{M_{2}}}{{\varepsilon}})$.
Recent work [KV25b, KV25a] revisited this problem by adapting $\mathsf{PS}$ to the general logconcave setting. Although $\mathsf{PS}$ can in principle be extended to any distribution where $\pi^{X|Y=y}\propto\pi^{X}(\cdot)\,\gamma_{h}(\cdot-y)$ can be sampled (e.g., when given access to a proximal oracle in the log-smooth setting), it is unclear how to implement this step for arbitrary $\pi^{X}$. To address this, [KV25b] introduced the exponential lifting technique: for $\pi^{X}\propto e^{-V}$, consider the lifted distribution
Here, the $X$-marginal of $\pi^{X,T}$ is exactly $\pi^{X}$, and $\mathcal{K}$ is convex whenever $\pi^{X}$ is logconcave. Thus, logconcave sampling reduces to the simpler problem of sampling from a logconcave distribution with a linear potential over a convex set. We then apply $\mathsf{PS}$ to this exponential distribution, denoted by $\mathsf{PS}_{\textup{exp}}$: for $z=(x,t)\in\mathbb{R}^{d+1}$ and $y\in\mathbb{R}^{d+1}$, it alternates $y\sim\pi^{Y|Z=z}=\mathcal{N}(z,hI_{d+1})$ and $z\sim\pi^{Z|Y=y}\propto\mathcal{N}(y-h\alpha,hI_{d+1})|_{\mathcal{K}}$ for $\alpha=de_{d+1}$, where the second step is implemented via rejection sampling, using the proposal $\mathcal{N}(y-h\alpha,hI_{d+1})$ with threshold $\tau$.
In [KV25b], it was shown that $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ and $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ with complexity $\widetilde{O}(M_{\infty}d^{2}(d\vee R^{2})\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. As in the uniform sampling case, the order of initial warmness was relaxed to $\mathcal{R}_{c}$ with $c=\widetilde{O}(1)$ in [KV25a], where $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ achieves query complexity $\widetilde{O}(M_{c}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. However, it remained open to balance the divergence order of initial warmness and final output distribution guarantee.
Extension to logconcave sampling (§5).
Our $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ with restart aligns the divergence order of initial warmness with that of the final error guarantee, from an $O(1)$-warm start, at an additive cost of $qd^{2}$ queries. This refines a long line of work on sampling in the classical setting. See §5.1 for further details.
Theorem 1.7 (Zeroth-order logconcave sampling from a warm start).
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, given by an evaluation oracle, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}\subset\mathsf{L}_{\pi,g}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$. Given $\varepsilon>0$, an initial distribution $\pi_{0}$, and $q\geq 2\vee\widetilde{\Omega}(\log\{d^{2}\,(\Lambda\vee 1)\log\frac{1}{\varepsilon}\})$ such that $M_{q}=\lVert\frac{\mathrm{d}\pi_{0}}{\mathrm{d}\pi}\rVert_{L^{q}(\pi)}\leq 10$, consider the algorithm that runs $\mathsf{PS}_{\textup{exp}}$ initialized at $\pi_{0}$ and restarts if $\mathsf{PS}_{\textup{exp}}$ fails before completing $N=\widetilde{O}(qd^{2}\,(\Lambda\vee 1)\log^{2}\frac{1}{\varepsilon})$ iterations. There exists choices of $h$ and $\tau$ such that this algorithm restarts with probability at most $1/50$, and its output $Z$ satisfies $\mathcal{R}_{q}(\operatorname{law}Z\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(qd^{2}\,(\Lambda\vee 1)\log^{3}\frac{1}{\varepsilon})$ evaluation queries in expectation.
We remark on a connection to the well-conditioned setting. In this case, the ground set contains a ball of radius $\Theta((d/\beta)^{1/2})$ centered at $\arg\min V$. Since $\Lambda\leq\alpha^{-1}$ by the Brascamp–Lieb or Lichnerowicz inequality, $\mathsf{PS}_{\textup{exp}}$ uses $\widetilde{O}(q\,(\kappa d+d^{2})\log\nicefrac{{1}}{{\varepsilon}})$ queries from an $O(1)$-warm start in $\mathcal{R}_{q}$. With the additional $qd^{2}$ queries, this matches the known $\widetilde{O}(q\kappa d\log\nicefrac{{1}}{{\varepsilon}})$-complexity results for zeroth-order samplers in this setting which additionally use a proximal oracle. See §B for further details.
Annealing approaches for uniform sampling have also been extended to arbitrary logconcave distributions [LV06a, KV25b, KV25a]. In the general case, the best-known complexity for warm-start generation is comparable to that of the convex-body case for $\mathsf{TV}$ and $\mathcal{R}_{\infty}$-warmness, with algorithms requiring additional $d^{2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$ queries. In §5.2, we extend our refined warm-start generation to general logconcave distributions.
Theorem 1.8.
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, given by an evaluation oracle, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. For any $\varepsilon_{\exp}\in(0,\frac{1}{10})$ and $q\geq 2\vee\widetilde{\Omega}(\log\frac{dD}{\varepsilon_{\exp}})$, there exists an annealing algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq 2$, using $\widetilde{O}(d^{2.5}+qd^{2}\,(R^{3/2}\vee 1)(\Lambda^{1/4}\vee 1))$ evaluation queries in expectation.
Combining the two theorems above, we establish the query complexity of sampling from an arbitrary logconcave distribution in $\mathbb{R}^{d}$ with $\mathcal{R}_{q}$-divergence guarantees, assuming only an evaluation oracle.
Corollary 1.9.
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, given by an evaluation oracle, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. For any $\varepsilon\in(0,\frac{1}{10})$ and $q\geq 1$, there exists an algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(d^{2.5}+qd^{2}R^{3/2}\,(\Lambda^{1/4}\vee 1)\log\frac{1}{\varepsilon}+qd^{2}\,(\Lambda\vee 1)\log^{4}\frac{1}{\varepsilon})$ evaluation queries in expectation.
1.1.4 Obstruction to further acceleration
A natural question is whether the annealing complexity in Corollary 1.9 can be improved beyond $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$. For isotropic logconcave distributions in particular, given the $d^{3}$ query complexity of [KV25b] for $\mathcal{R}_{\infty}$-guarantees and our $qd^{2.75}$ complexity for $\mathcal{R}_{q}$-guarantees, it is natural to ask whether the $d^{2.5}$ iteration complexity of the $\mathsf{Speedy\ walk}$ established in [LV24] can be matched as a query complexity, at least for $\mathsf{TV}$-guarantees (note that the speedy walk is an abstraction that, in general, does not have an efficient implementation).
Towards this goal, [KV25a] conjectured that for the uniform distribution over any isotropic convex body $\pi$ in $\mathbb{R}^{d}$, the Gaussian-tilted distribution $\pi\gamma_{t}$ satisfies $\lVert\operatorname{cov}\pi\gamma_{t}\rVert=\widetilde{O}(1)$ for all $t>0$, a property called $\widetilde{O}(1)$-quadratic-tilt stability. Under this conjecture, one can afford more aggressive annealing steps of size $\sigma_{i+1}^{2}=\sigma_{i}^{2}(1+\sigma_{i}^{2}/R)$ (instead of $\sigma_{i+1}^{2}=\sigma_{i}^{2}(1+\sigma_{i}/R)$), leading to an improved warm-start generation complexity of $\widetilde{O}(d^{2.5})$ for isotropic logconcave distributions (see §6.1). We verify in §A that structured bodies—including the hypercube, the simplex, and any isotropic convex body of revolution—satisfy $O(1)$-quadratic-tilt stability, confirming the desired $\widetilde{O}(d^{2.5})$ query complexity of sampling from these structured distributions.
However, in §6.2 we disprove the conjecture in general, establishing a fundamental obstruction to this line of acceleration. Specifically, adapting a construction of Bizeul [BIZ26], we show that there exists an isotropic convex body whose Gaussian-tilted distribution exhibits a peak variance of $d^{1/3}$ (Proposition 6.3), disproving the conjecture and proving a lower bound for the Gaussian annealing approach that has been central to this line of work since [CV14, CV15, CV18].
@@ 1.2 Preliminaries
We call a function $f:\mathbb{R}^{d}\to[0,\infty)$ logconcave if $\log f$ is convex in $\mathbb{R}^{d}$, and a probability measure $\pi$ (or distribution) logconcave if it has a logconcave density function with respect to the Lebesgue measure. We abuse notation by using the same symbol for a distribution and density. We assume that any logconcave distributions considered in this work are non-degenerate; otherwise, we could simply work on an affine subspace on which degenerate distributions are supported. For $t\geq 0$, $\pi$ is called $t$-strongly logconcave if $\log\pi$ is $t$-strongly convex (i.e., $\log\pi-\frac{t}{2}\,|\cdot|^{2}$ is convex). Clearly, log-concavity is preserved under multiplication, and a classical result by Prékopa and Leindler ensures that the convolution also preserves log-concavity. A distribution is called isotropic if its barycenter is at the origin and has the identity covariance matrix. Note that as the logconcave distributions decay exponentially fast at infinity, they have finite moments of all orders.
@@
For $\lambda=q/(2q-1)$, we have $\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\frac{2q}{2q-1}\,\mathcal{R}_{2q-1}(\mu\mathbin{\|}\nu)+\mathcal{R}_{2q}(\nu\mathbin{\|}\pi)$. We refer readers to [vH14, MIR17] for more properties of the Rényi divergence.
2 Uniform sampling from convex bodies with balanced Rényi divergence
We begin by describing our approach for the general proximal sampler ($\mathsf{PS}$) and then focus specifically on $\mathsf{PS}_{\textup{unif}}$ for uniform sampling in §2.2, which was proposed and analyzed in [KVZ24]. $\mathsf{PS}$ with step size $h$ alternates between two steps: starting with $k=0$ and $x_{0}\sim\pi_{0}^{X}$, (i) sample $y_{k+1}\sim\pi^{Y|X=x_{k}}$ and (ii) sample $x_{k+1}\sim\pi^{X|Y=y_{k+1}}$. Precisely, let $P_{h}$ and $Q_{h}$ be the transition kernels of the forward and backward steps, respectively, defined as
@@ 2 Uniform sampling from convex bodies with balanced Rényi divergence
Assume that there exists a suitable proposal distribution $\mu_{y}$ such that $(\mu_{y}\mid\text{accepted})=Q_{h}(y,\cdot)$. For uniform sampling, we take $\mu_{y}=\mathcal{N}(y,hI_{d})$ and accept the proposal if it lies inside $\mathcal{K}$. Given this setup, we implement the backward step via rejection sampling with proposal $\mu_{y_{k+1}}$. If the rejection loop exceeds $\tau$ trials, we declare failure and restart from scratch with a new sample from $\pi_{0}^{X}$.
Kook and Vempala [KV25a] obtained the query complexity of $\mathsf{PS}$ for obtaining a uniform sample whose law is $\varepsilon$-close to the target $\pi^{X}$ in $\mathcal{R}_{2}$-divergence, assuming the warmness is given in $\mathcal{R}_{q}$ with $q\geq\widetilde{O}(1)$. We refine their analysis to obtain the query complexity of $\mathsf{PS}_{\textup{unif}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ given $\lVert\mathrm{d}\pi_{0}^{X}/\mathrm{d}\pi^{X}\rVert_{L^{q}(\pi^{X})}\leq 10$ (i.e., an $O(1)$-warm start in $\mathcal{R}_{q}$). To this end, we decouple the analysis into two components: (1) a mixing analysis and (2) the query complexity of the backward step. The first part determines how many iterations $N$ are required for the sample $X_{N}$ to be $\varepsilon$-close to $\pi^{X}$. The second determines how many queries $\mathsf{PS}$ uses in each iteration. Therefore, the total query complexity of the algorithm is the product of these two components.